OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (triciaal):

one approach (the long way) 5>2 f(5) = f(1) + f(2) + f(5 - 1) f(4) = (f1) + f(2) + f(4 - 1) f(3) = f(1) + f(2) + f(3-1) f(5)= f(1) +f(2) + f(1) + f(2) + f(1) + f(2) + f(2) f(5) = 3*f(1) + 3f(2) = 3*(2 + 5) = 21

OpenStudy (triciaal):

correction I missed an (f2) should be 3*f(1) + 4*f(2) so it would be 21 + 5 = 26